# The interplay of WGC and WCCC via charged scalar field fluxes in the RPST framework

### v1

@@ 4 Charged Scalar Field and RPS Thermodynamics
In RPS thermodynamics, the radius AdS is fixed, and Newton’s constant is considered as a variable. It can also be found that by attracting a scalar field or a particle into the black hole, there are changes in the quantities $(M,\hat{Q},C,r_{h})$ of the black hole, which, according to relations (6) and (11), also leads to changes in the metric of the black hole. Also, WCCC, the event horizon of the black hole plays an important role and is obtained from the relation $h(M,\hat{Q},r_{h},C)=0$. Therefore, in the case where the black hole attracts the scalar field or particle, this event horizon must exist, in this case, the event horizon of the black hole changes and is obtained from the following relationship,
$h(M+dM,\hat{Q}+d\hat{Q},r_{h}+dr_{h},C+dC)=0,$ (31)
@@
$\begin{split}&h(M+dM,\hat{Q}+d\hat{Q},r_{h}+dr_{h},C+dC)=h(M,\hat{Q},r_{h},C)\\ &\qquad+\frac{\partial h}{\partial M}|_{r=r_{h}}dM+\frac{\partial h}{\partial\hat{Q}}|_{r=r_{h}}d\hat{Q}+\frac{\partial h}{\partial r}|_{r=r_{h}}dr+\frac{\partial h}{\partial C}|_{r=r_{h}}dC\end{split}$ (32)
@@
$\begin{split}&h(M,\hat{Q},r_{h},C)=0,\qquad\frac{\partial h}{\partial M}|_{r=r_{h}}=-\frac{2\ell^{2}}{Cr_{h}},\qquad\frac{\partial h}{\partial\hat{Q}}|_{r=r_{h}}=\frac{2\ell^{2}\hat{Q}}{C^{2}r_{h}^{2}},\\ &\frac{\partial h}{\partial r}|_{r=r_{h}}=-\frac{2\ell^{2}\hat{Q}^{2}}{C^{2}r_{h}^{3}}+\frac{2\ell^{2}M}{Cr_{h}^{2}}+\frac{2r_{h}}{\ell^{2}}=4\pi T,\qquad\frac{\partial h}{\partial C}|_{r=r_{h}}=\frac{1}{C}\bigg{(}\frac{2\ell^{2}M}{Cr_{h}}-\frac{2\ell^{2}\hat{Q}^{2}}{C^{2}r_{h}^{2}}\bigg{)}\end{split}$ (33)
By using equations (28), (29), (30), (31), (31) and (33), one can obtain the variations in the event horizon,
$\begin{split}dr_{h}=\frac{\ell^{2}r_{h}(\omega-\hat{q}\hat{\Phi})^{2}}{2\pi CT}dt-\frac{1}{4\pi CT}\bigg{(}1+\frac{r_{h}^{2}}{\ell^{2}}\bigg{)}dC\end{split}$ (34)
Using equation (12), we calculate the variations of entropy,
$\begin{split}dS=\frac{2\pi Cr_{h}}{\ell^{2}}dr_{h}+\frac{\pi r_{h}^{2}}{\ell^{2}}dC\end{split}$ (35)
By combining equations (34) and (35), we rewrite the variations of entropy ,
$\begin{split}dS=\frac{r_{h}^{2}(\omega-\hat{q}\hat{\Phi})^{2}}{T}dt-\frac{r_{h}}{2\ell^{2}T}\bigg{(}1+\frac{r_{h}^{2}}{\ell^{2}}-2\pi r_{h}T\bigg{)}dC\end{split}$ (36)
@@
$\begin{split}T=\frac{1}{2\pi r_{h}}\bigg{(}1+\frac{r_{h}^{2}}{\ell^{2}}\bigg{)}\end{split}$ (37)
despite maintaining the conditions of the RPS, the entropy changes will still reduce to the solution of the normal space. According to equation (36), we find that entropy changes can be negative; in this case, Newton’s second law can be violated. If the Hawking temperature of the black hole has a lower bound, the entropy changes can always be positive,
$\begin{split}T\geq\frac{1}{2\pi r_{h}}\bigg{(}1+\frac{r_{h}^{2}}{\ell^{2}}\bigg{)}.\end{split}$ (38)
@@
5 WCCC in Near-Extremal and Extremal Charged AdS black hole
The WCCC suggests that a black hole’s singularity should be hidden by the horizon from an observer at infinity, preventing a naked singularity within the black hole’s structure if it has a stable horizon. Our current study explores the stability of the outer horizon during the scattering of a charged scalar field, considering RPS thermodynamic. By absorbing the scalar field, the initial state $h(M,Q,r_{h},C)$ transitions to the final state $h(M+dM,Q+dQ,r_{h}+dr_{h},C+dC)$ over a small time interval $dt$. By setting the outer horizon as $h(M,Q,r,C)=0$, we can determine the presence of the horizon by examining solutions in $h(M+dM,Q+dQ,r_{h}+dr_{h},C+dC)=0$. This analysis simplifies to studying the change in the minimum value of $h$.
@@ 5 WCCC in Near-Extremal and Extremal Charged AdS black hole
$\begin{split}h_{min}\equiv h(r_{min})=1-\frac{2\ell^{2}M}{Cr_{min}}+\frac{\ell^{2}\hat{Q}^{2}}{C^{2}r_{min}^{2}}+\frac{r_{min}^{2}}{\ell^{2}}\end{split}$ (39)
$\begin{split}\frac{\partial h_{min}}{\partial r_{min}}=-\frac{2\ell^{2}\hat{Q}^{2}}{C^{2}r_{min}^{3}}+\frac{2\ell^{2}M}{Cr_{min}^{2}}+\frac{2r_{min}}{\ell^{2}}=0\end{split}$ (40)
$\begin{split}\frac{\partial^{2}h_{min}}{\partial r_{min}^{2}}=\frac{6\ell^{2}\hat{Q}^{2}}{C^{2}r_{min}^{4}}-\frac{4\ell^{2}M}{Cr_{min}^{3}}+\frac{2}{\ell^{2}}>0.\end{split}$ (41)
@@
$\begin{split}h(M+dM,\hat{Q}+d\hat{Q},C+dC,r_{min}+dr_{min})=h_{min}+\frac{\partial h_{min}}{\partial M}dM+\frac{\partial h_{min}}{\partial\hat{Q}}d\hat{Q}+\frac{\partial h_{min}}{\partial C}dC+\frac{\partial h_{min}}{\partial r_{min}}dr_{min}\end{split}$ (42)
@@
$\begin{split}&h_{min}=\delta\leq 0,\qquad\frac{\partial h_{min}}{\partial M}=-\frac{2\ell^{2}}{Cr_{min}},\qquad\frac{\partial h_{min}}{\partial\hat{Q}}=\frac{2\ell^{2}\hat{Q}}{C^{2}r_{min}^{2}},\\ &\frac{\partial h_{min}}{\partial C}=\frac{1}{C}\bigg{(}\frac{2\ell^{2}M}{Cr_{min}}-\frac{2\ell^{2}\hat{Q}^{2}}{C^{2}r_{min}^{2}}\bigg{)},\qquad\frac{\partial h_{min}}{\partial r_{min}}=\frac{C^{2}r_{h}r_{min}\left(\ell^{2}r_{h}+r_{h}^{3}+2r_{min}^{3}\right)-\ell^{4}\hat{Q}^{2}\left(2r_{h}-r_{min}\right)}{C^{2}\ell^{2}r_{h}r_{min}^{3}}=0\end{split}$ (43)
By putting relations (28), (29), (30) and (34) in relation (42), we get,
$\begin{split}&h(M+dM,\hat{Q}+d\hat{Q},C+dC,r_{min}+dr_{min})=\\ &-\frac{2\ell^{2}r_{h}(\omega-\hat{q}\hat{\phi}_{h})\big{[}r_{h}(\omega-\hat{q}\hat{\phi}_{min})-r_{min}(\omega-\hat{q}\hat{\phi}_{h})\big{]}}{Cr_{min}}dt+\frac{\left(r_{h}-r_{min}\right)\left[C^{2}r_{h}r_{min}\left(\ell^{2}+r_{h}^{2}\right)-\ell^{4}\hat{Q}^{2}\right]}{C^{3}\ell^{2}r_{h}r_{min}^{2}}dC\end{split}$ (44)
@@
$\begin{split}h(M+dM,\hat{Q}+d\hat{Q},C+dC,r_{min}+dr_{min})=0\end{split}$ (45)
Thus, despite the scattering, the black hole remains in an extremal state, and the WCCC is valid. Also, the result obtained in the RPS is the same as the EPS and the normal space. To study the near-extremal black hole, we analyze the event horizon close to the minimum point of $f(r_{min})$, i.e. at $r_{h}=r_{min}+\epsilon$ and $10<\epsilon<<$ 1. In this scenario, equation (44) can be expressed as follows,
$\begin{split}&h(M+dM,\hat{Q}+d\hat{Q},C+dC,r_{min}+dr_{min})=\delta-\epsilon\left[\frac{2\ell^{2}(\omega-\hat{q}\hat{\phi}_{h})(\omega-2\hat{q}\hat{\phi}_{h})}{C}+\mathcal{O}(\epsilon)\right]dt\\ &+\epsilon\left[\frac{C^{2}r_{h}^{2}\left(\ell^{2}+r_{h}^{2}\right)-\ell^{4}\hat{Q}^{2}}{C^{3}\ell^{2}r_{h}^{3}}+\mathcal{O}(\epsilon)\right]dC\end{split}$ (46)
@@
$\begin{split}h_{min}=\delta=-\epsilon\bigg{[}\frac{C^{2}\ell^{2}r_{h}^{2}+3C^{2}r_{h}^{4}-\ell^{4}\hat{Q}^{2}}{C^{2}\ell^{2}r_{h}^{3}}+\mathcal{O}(\epsilon)\bigg{]}\end{split}$ (47)
Therefore, for small time changes that lead to small charge center changes, i.e. $dt\sim\epsilon$ and $dC\sim\epsilon$, equation (46) is obtained as follows,
$\begin{split}h(M+dM,\hat{Q}+d\hat{Q},C+dC,r_{min}+dr_{min})=\delta\pm\mathcal{O}(\epsilon^{2})<0\end{split}$ (48)
Hence, the adjustment to the minimum value can be considered negligible in the initial order of d. Consequently, it can be inferred that there is no alteration to the condition of the black hole: a nearly extremal black hole remains nearly extremal even with variations in mass and electric charge. Thus, the WCCC is valid for near-extremal black holes as well. However, according to relations (46) and (47), we find that, in the condition that $dC>>1$, there is a possibility that the conjecture of weak cosmic censorship is violated. Also, when there are no changes in the central charge, the result obtained is in agreement with the result obtained in the extended phase space [36].
@@ 6 WGC with equivalence of mass and energy
$\begin{split}E=\mu_{s},\qquad E=\omega\qquad\Rightarrow\qquad\mu_{s}=\omega\end{split}$ (49)
Also, according to relation (7), we can have,
$q\Phi_{h}=\hat{q}\hat{\Phi}_{h}$
@@
Table 1: near-extremal (N- E) , slightly extremal (S- E)
Using the equivalence of mass and energy and relation (46), we find that the black hole emits particles in its radiation that obey the weak gravity conjecture, which nevertheless holds the weak cosmic censorship conjecture, so that the black hole from The extremity becomes more distant. But in the case where the black hole attracts the particle, two situations occur, in the first case, the black hole attracts the particle that follows the WGC. In this case, the WCCC is valid, with the difference that the black hole is closer to the extremity state. In the second case, the black hole attracts a particle that does not follow the WGC, the WCCC will be valid, so that the black hole moves away from the extremity state. Next, we obtain the electric potential of the black hole near the extreme state. Given $r_{h}=r_{min}+\epsilon$, the electric potential can be expressed as,
$\begin{split}\Phi_{h}=\frac{Q}{r_{h}}=\frac{Q}{r_{min}+\epsilon}=\frac{\hat{Q}}{r_{min}\sqrt{C}}(1-\frac{\epsilon}{r_{min}})\end{split}$ (50)
@@
$\begin{split}r_{min}=\frac{\sqrt{G}}{\sqrt{6}}\bigg{(}\sqrt{C^{2}+12\hat{Q}^{2}}-C\bigg{)}^{\frac{1}{2}}\end{split}$ (51)
By placing (50) and (51) into Superradiance conditions ( $\frac{1}{\Phi_{h}}<\frac{q}{\mu_{s}}$), we have,
$\begin{split}\frac{q}{\mu_{s}}>\sqrt{G}\big{(}1+\frac{\epsilon}{r_{min}}\big{)}\frac{\sqrt{C}\big{(}\sqrt{C^{2}+12\hat{Q}^{2}}-C\big{)}^{\frac{1}{2}}}{\sqrt{6}\hat{Q}}\end{split}$ (52)
According to relation (52), in the condition that $C>>\hat{Q}$, we have $\frac{q}{\mu_{s}}>\sqrt{G}\big{(}1+\frac{\epsilon}{r_{min}}\big{)}$, in which case the particles emitted from the black hole follow the WGC. But in the case where the black hole absorbs the particle, even when $C>>\hat{Q}$, we have,
$\begin{split}\frac{\sqrt{G}}{2}\big{(}1+\frac{\epsilon}{r_{min}}\big{)}<\frac{q}{\mu_{s}}<\sqrt{G}\big{(}1+\frac{\epsilon}{r_{min}}\big{)}\end{split}$ (53)
Therefore, the WGC does not hold in this case. The black hole attracts particles that do not obey the WGC and repels particles that obey the WGC. In both cases the WCCC holds. Also, according to relation (53), we find that only if the black hole has a very small event horizon and is in an extreme state $(r_{min}=\epsilon)$ can absorb particles that follow the WGC. Therefore, these small black holes get closer to their extreme limit, which increases the probability of the black hole collapsing.

### v2

@@ 4 Charged scalar field and RPS thermodynamics
In RPS thermodynamics, the radius AdS is fixed, and Newton’s constant is considered as a variable. It can also be found that by attracting a scalar field or a particle into the black hole, there are changes in the quantities $(M,\hat{Q},C,r_{h})$ of the black hole, which, according to relations (2.6) and (2.11), also leads to changes in the metric of the black hole. Also, WCCC, the event horizon of the black hole plays an important role and is obtained from the relation $h(M,\hat{Q},r_{h},C)=0$. Therefore, in the case where the black hole attracts the scalar field or particle, this event horizon must exist, in this case, the event horizon of the black hole changes and is obtained from the following relationship,
$h(M+dM,\hat{Q}+d\hat{Q},r_{h}+dr_{h},C+dC)=0,$ (4.1)
@@
$\begin{split}&h(M+dM,\hat{Q}+d\hat{Q},r_{h}+dr_{h},C+dC)=h(M,\hat{Q},r_{h},C)\\ &\qquad+\frac{\partial h}{\partial M}|_{r=r_{h}}dM+\frac{\partial h}{\partial\hat{Q}}|_{r=r_{h}}d\hat{Q}+\frac{\partial h}{\partial r}|_{r=r_{h}}dr+\frac{\partial h}{\partial C}|_{r=r_{h}}dC\end{split}$ (4.2)
@@
$\begin{split}&h(M,\hat{Q},r_{h},C)=0,\qquad\frac{\partial h}{\partial M}|_{r=r_{h}}=-\frac{2\ell^{2}}{Cr_{h}},\qquad\frac{\partial h}{\partial\hat{Q}}|_{r=r_{h}}=\frac{2\ell^{2}\hat{Q}}{C^{2}r_{h}^{2}},\\ &\frac{\partial h}{\partial r}|_{r=r_{h}}=-\frac{2\ell^{2}\hat{Q}^{2}}{C^{2}r_{h}^{3}}+\frac{2\ell^{2}M}{Cr_{h}^{2}}+\frac{2r_{h}}{\ell^{2}}=4\pi T,\qquad\frac{\partial h}{\partial C}|_{r=r_{h}}=\frac{1}{C}\bigg{(}\frac{2\ell^{2}M}{Cr_{h}}-\frac{2\ell^{2}\hat{Q}^{2}}{C^{2}r_{h}^{2}}\bigg{)}\end{split}$ (4.3)
By using equations (3.13), (3.14), (3.15), (4.1), (4.1) and (4.3), one can obtain the variations in the event horizon,
$\begin{split}dr_{h}=\frac{\ell^{2}r_{h}(\omega-\hat{q}\hat{\Phi})^{2}}{2\pi CT}dt-\frac{1}{4\pi CT}\bigg{(}1+\frac{r_{h}^{2}}{\ell^{2}}\bigg{)}dC\end{split}$ (4.4)
Using equation (2.12), we calculate the variations of entropy,
$\begin{split}dS=\frac{2\pi Cr_{h}}{\ell^{2}}dr_{h}+\frac{\pi r_{h}^{2}}{\ell^{2}}dC\end{split}$ (4.5)
By combining equations (4.4) and (4.5), we rewrite the variations of entropy ,
$\begin{split}dS=\frac{r_{h}^{2}(\omega-\hat{q}\hat{\Phi})^{2}}{T}dt-\frac{r_{h}}{2\ell^{2}T}\bigg{(}1+\frac{r_{h}^{2}}{\ell^{2}}-2\pi r_{h}T\bigg{)}dC.\end{split}$ (4.6)
@@
$\begin{split}T=\frac{1}{2\pi r_{h}}\bigg{(}1+\frac{r_{h}^{2}}{\ell^{2}}\bigg{)}.\end{split}$ (4.7)
Despite maintaining the conditions of the RPS, the entropy changes will still reduce to the solution of the normal space. According to equation (4.6), we find that entropy changes can be negative; in this case, Newton’s second law can be violated. If the Hawking temperature of the black hole has a lower bound, the entropy changes can always be positive,
$\begin{split}T\geq\frac{1}{2\pi r_{h}}\bigg{(}1+\frac{r_{h}^{2}}{\ell^{2}}\bigg{)}.\end{split}$ (4.8)
@@
5 WCCC in near-extremal and extremal cases
The WCCC suggests that a black hole’s singularity should be hidden by the horizon from an observer at infinity, preventing a naked singularity within the black hole’s structure if it has a stable horizon. Our current study explores the stability of the outer horizon during the scattering of a charged scalar field, considering RPS thermodynamics. By absorbing the scalar field, the initial state $h(M,Q,r_{h},C)$ transitions to the final state $h(M+dM,Q+dQ,r_{h}+dr_{h},C+dC)$ over a small time interval $dt$. By setting the outer horizon as $h(M,Q,r,C)=0$, we can determine the presence of the horizon by examining solutions in $h(M+dM,Q+dQ,r_{h}+dr_{h},C+dC)=0$. This analysis simplifies studying the change in the minimum value of $h$.
@@ 5 WCCC in near-extremal and extremal cases
$\begin{split}h_{min}\equiv h(r_{min})=1-\frac{2\ell^{2}M}{Cr_{min}}+\frac{\ell^{2}\hat{Q}^{2}}{C^{2}r_{min}^{2}}+\frac{r_{min}^{2}}{\ell^{2}},\end{split}$ (5.1)
$\begin{split}\frac{\partial h_{min}}{\partial r_{min}}=-\frac{2\ell^{2}\hat{Q}^{2}}{C^{2}r_{min}^{3}}+\frac{2\ell^{2}M}{Cr_{min}^{2}}+\frac{2r_{min}}{\ell^{2}}=0,\end{split}$ (5.2)
$\begin{split}\frac{\partial^{2}h_{min}}{\partial r_{min}^{2}}=\frac{6\ell^{2}\hat{Q}^{2}}{C^{2}r_{min}^{4}}-\frac{4\ell^{2}M}{Cr_{min}^{3}}+\frac{2}{\ell^{2}}>0.\end{split}$ (5.3)
@@
$\begin{split}h(M+dM,\hat{Q}+d\hat{Q},C+dC,r_{min}+dr_{min})=h_{min}+\frac{\partial h_{min}}{\partial M}dM+\frac{\partial h_{min}}{\partial\hat{Q}}d\hat{Q}+\frac{\partial h_{min}}{\partial C}dC+\frac{\partial h_{min}}{\partial r_{min}}dr_{min},\end{split}$ (5.4)
@@
$\begin{split}&h_{min}=\delta\leq 0,\qquad\frac{\partial h_{min}}{\partial M}=-\frac{2\ell^{2}}{Cr_{min}},\qquad\frac{\partial h_{min}}{\partial\hat{Q}}=\frac{2\ell^{2}\hat{Q}}{C^{2}r_{min}^{2}},\\ &\frac{\partial h_{min}}{\partial C}=\frac{1}{C}\bigg{(}\frac{2\ell^{2}M}{Cr_{min}}-\frac{2\ell^{2}\hat{Q}^{2}}{C^{2}r_{min}^{2}}\bigg{)},\qquad\frac{\partial h_{min}}{\partial r_{min}}=\frac{C^{2}r_{h}r_{min}\left(\ell^{2}r_{h}+r_{h}^{3}+2r_{min}^{3}\right)-\ell^{4}\hat{Q}^{2}\left(2r_{h}-r_{min}\right)}{C^{2}\ell^{2}r_{h}r_{min}^{3}}=0.\end{split}$ (5.5)
By putting relations (3.13), (3.14), (3.15) and (4.4) in relation (5.4), we get,
$\begin{split}&h(M+dM,\hat{Q}+d\hat{Q},C+dC,r_{min}+dr_{min})=\\ &-\frac{2\ell^{2}r_{h}(\omega-\hat{q}\hat{\phi}_{h})\big{[}r_{h}(\omega-\hat{q}\hat{\phi}_{min})-r_{min}(\omega-\hat{q}\hat{\phi}_{h})\big{]}}{Cr_{min}}dt+\frac{\left(r_{h}-r_{min}\right)\left[C^{2}r_{h}r_{min}\left(\ell^{2}+r_{h}^{2}\right)-\ell^{4}\hat{Q}^{2}\right]}{C^{3}\ell^{2}r_{h}r_{min}^{2}}dC.\end{split}$ (5.6)
@@
$\begin{split}h(M+dM,\hat{Q}+d\hat{Q},C+dC,r_{min}+dr_{min})=0.\end{split}$ (5.7)
Thus, despite the scattering, the black hole remains in an extremal state, and the WCCC is valid. Also, the result obtained in the RPS is the same as the EPS and the normal space. To study the near-extremal black hole, we analyze the event horizon close to the minimum point of $f(r_{min})$, i.e. at $r_{h}=r_{min}+\epsilon$ and $10<\epsilon<<$ 1. In this scenario, equation (5.6) can be expressed as follows,
$\begin{split}&h(M+dM,\hat{Q}+d\hat{Q},C+dC,r_{min}+dr_{min})=\delta-\epsilon\left[\frac{2\ell^{2}(\omega-\hat{q}\hat{\phi}_{h})(\omega-2\hat{q}\hat{\phi}_{h})}{C}+\mathcal{O}(\epsilon)\right]dt\\ &+\epsilon\left[\frac{C^{2}r_{h}^{2}\left(\ell^{2}+r_{h}^{2}\right)-\ell^{4}\hat{Q}^{2}}{C^{3}\ell^{2}r_{h}^{3}}+\mathcal{O}(\epsilon)\right]dC.\end{split}$ (5.8)
@@
$\begin{split}h_{min}=\delta=-\epsilon\bigg{[}\frac{C^{2}\ell^{2}r_{h}^{2}+3C^{2}r_{h}^{4}-\ell^{4}\hat{Q}^{2}}{C^{2}\ell^{2}r_{h}^{3}}+\mathcal{O}(\epsilon)\bigg{]}\end{split}$ (5.9)
Therefore, for small time changes that lead to small charge center changes, i.e. $dt\sim\epsilon$ and $dC\sim\epsilon$, equation (5.8) is obtained as follows,
$\begin{split}h(M+dM,\hat{Q}+d\hat{Q},C+dC,r_{min}+dr_{min})=\delta\pm\mathcal{O}(\epsilon^{2})<0.\end{split}$ (5.10)
Hence, the adjustment to the minimum value can be considered negligible in the initial order of d. Consequently, it can be inferred that there is no alteration to the condition of the black hole: a nearly extremal black hole remains nearly extremal even with variations in mass and electric charge. Thus, the WCCC is valid for near-extremal black holes as well. However, according to relations (5.8) and (5.9), we find that, in the condition that $dC>>1$, there is a possibility that the conjecture of weak cosmic censorship is violated. Also, when there are no changes in the central charge, the result obtained is in agreement with the result obtained in the extended phase space [76].
@@ 6 WGC with equivalence of mass and energy
$\begin{split}E=\mu_{s},\qquad E=\omega\qquad\Rightarrow\qquad\mu_{s}=\omega.\end{split}$ (6.1)
Also, according to relation (2.7), we can have,
$q\Phi_{h}=\hat{q}\hat{\Phi}_{h}.$
@@
Table 1: Summary of the results
Using the equivalence of mass and energy and relation (5.8), we find that the black hole emits particles in its radiation that obey the weak gravity conjecture, which nevertheless holds the weak cosmic censorship conjecture so that the black hole from The extremity becomes more distant. But in the case where the black hole attracts the particle, two situations occur, in the first case, the black hole attracts the particle that follows the WGC. In this case, the WCCC is valid, with the difference that the black hole is closer to the extremity state. In the second case, the black hole attracts a particle that does not follow the WGC, the WCCC will be valid so that the black hole moves away from the extremity state. Next, we obtain the electric potential of the black hole near the extreme state. Given $r_{h}=r_{min}+\epsilon$, the electric potential can be expressed as,
$\begin{split}\Phi_{h}=\frac{Q}{r_{h}}=\frac{Q}{r_{min}+\epsilon}=\frac{\hat{Q}}{r_{min}\sqrt{C}}(1-\frac{\epsilon}{r_{min}}).\end{split}$ (6.2)
@@
$\begin{split}r_{min}=\frac{\sqrt{G}}{\sqrt{6}}\bigg{(}\sqrt{C^{2}+12\hat{Q}^{2}}-C\bigg{)}^{\frac{1}{2}}.\end{split}$ (6.3)
By placing (6.2) and (6.3) into Superradiance conditions ( $\frac{1}{\Phi_{h}}<\frac{q}{\mu_{s}}$), we have,
$\begin{split}\frac{q}{\mu_{s}}>\sqrt{G}\big{(}1+\frac{\epsilon}{r_{min}}\big{)}\frac{\sqrt{C}\big{(}\sqrt{C^{2}+12\hat{Q}^{2}}-C\big{)}^{\frac{1}{2}}}{\sqrt{6}\hat{Q}}.\end{split}$ (6.4)
According to relation (6.4), in the condition that $C>>\hat{Q}$, we have $\frac{q}{\mu_{s}}>\sqrt{G}\big{(}1+\frac{\epsilon}{r_{min}}\big{)}$, in which case the particles emitted from the black hole follow the WGC. But in the case where the black hole absorbs the particle, even when $C>>\hat{Q}$, we have,
$\begin{split}\frac{\sqrt{G}}{2}\big{(}1+\frac{\epsilon}{r_{min}}\big{)}<\frac{q}{\mu_{s}}<\sqrt{G}\big{(}1+\frac{\epsilon}{r_{min}}\big{)}.\end{split}$ (6.5)
Therefore, the WGC does not hold in this case. The black hole attracts particles that do not obey the WGC and repels particles that obey the WGC. In both cases the WCCC holds. Also, according to relation (6.5), we find that only if the black hole has a very small event horizon and is in an extreme state $(r_{min}=\epsilon)$ can absorb particles that follow the WGC. Therefore, these small black holes get closer to their extreme limit, which increases the probability of the black hole collapsing.